# Mixed Precision Iterative Refinement for Linear Inverse Problems

### v2

@@ 3              $\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\in\mathbb{R},\quad\text{for }j=1,2,\dots,n$ (in Pr1)
4  Output: $\Psi^{(k+1)}\in\mathbb{R}^{n}$
5  for $k=0,1,2,\dots$ do
@@ 4  Output: $\Psi^{(k+1)}\in\mathbb{R}^{n}$
7              $q_{j}^{(k)}=\sigma_{A,j}^{2}\left(1-\sigma^{2}_{A,j}\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}\right)^{k}$ $[\text{in }\text{Pr}_{3}]$
8              $g_{j}^{(k)}=\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}q_{j}^{(k)}$ $[\text{in }\text{Pr}_{3}]$
9              $\psi_{j}^{(k+1)}=\psi_{j}^{(k)}+g_{j}^{(k)}$ $[\text{in }\text{Pr}_{3}]\,\,$
@@
with the $j^{th}$ filter factor per $k^{th}$ iterative step with indicated precision given by
@@
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $\Phi^{(0)}=0\in\mathbb{R}^{n},\,\,\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\alpha>0$
3  Output: $\Phi^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\in\mathbb{R},\,j=1,\dots,n$ from $M^{T}M=\left(A^{T}A+\alpha^{2}I\right)$ $[\text{in }\text{Pr}_{1}]$
5  for $k=0,1,2,\dots$ do
6        Compute $\Psi^{(k+1)}$ using Algorithm 4 $[\text{in }\text{Pr}_{3}]$
@@
8              $q_{j}^{(k)}=\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\left[\psi_{j}^{(k+1)}-\psi_{j}^{(k)}\right]-\alpha^{2}\phi^{(k-1)}_{j}$ $[\text{in }\text{Pr}_{3}]$
9              $q_{j}^{(k)}=q_{j}^{(k)}+\left(\frac{\alpha^{2}\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-2}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-2-i)}$ $[\text{in }\text{Pr}_{3}]$
10              $g_{j}^{(k)}=\left(\sigma_{M,j}^{2}+\alpha^{2}\right)^{-1}q_{j}^{(k)}$ $[\text{in }\text{Pr}_{2}]$
11              $\phi_{j}^{(k+1)}=\phi_{j}^{(k)}+g_{j}^{(k)}$ $[\text{in }\text{Pr}_{2}]$
@@ 5.1 Effective filter factors
To assess the experimental veracity of our filter factor results from Section 4 for IR on the Tikhonov problem we compare against experimentally derived filter factors which we term effective filter factors. Using (5), the effective filter factors $\omega^{(k)}_{j}$ of an approximate solution at the $k^{th}$ iteration may be computed as follows
@@ 5.2 Low precision simulation in MATLAB
To ascertain the effectiveness of mixed precision IR on the Tikhonov problem and the experimental validity of the results presented in Section 4 we utilize the software package chop introduced by Higham and Pranesh in [18]. The chop function simulates lower precision arithmetic by rounding array entries given in a MATLAB native precision (e.g. single or double) to a target precision which is stored in a higher precision with non-utilized representation bits set to zero. Chen et al. in [6] consider the application of chop for the solution of structured inverse problems and investigate associated numerical considerations of its usage.
Various target precisions are supported by the chop software including user-customizable ones. The precisions we utilize in our numerical experiments in Section 5.3 and their associated shorthand notations are provided in Table 1. We comment that we did not experience any significant experimental differences when using precisions that did not support subnormal numbers, e.g. Google’s bfloat16111https://en.wikipedia.org/wiki/Bfloat16__\__floating-point__\__format, compared to those due to the IEEE standard [^2] https://en.wikipedia.org/wiki/IEEE__\__754##\##2019 during our numerical investigations and therefore do not include them in our results.
@@
Pr Shorthand Name Exponent Bits Mantissa Bits Approx. Decimal Round-off IEEE Standard
$1$ fp64 $11$ $52$ $-161.11e\text{-}$ 16 Yes
$2$ fp32 $8$ $23$ $-85.96e\text{-}$ 8 Yes
$3$ fp16 $5$ $10$ $-44.88e\text{-}$ 4 Yes
@@ 5.3 Examples and Results
To realistically simulate the inverse problem, noise was added to the true right-hand side, $\hat{b}$, by forming the vector $e$ with normally distributed random entries with mean zero so that $b=Ax+e$; the vector $e$ is scaled so as to correspond to a specific noise level given by $\mu=100\left(\|e\|/\|\hat{b}\|\right)$. We will refer to $\mu$ as the noise level. The condition number of the matrix $A$ as determined by the MATLAB function cond() is $\approx 10^{9}$; it can also be easily verified that the singular values of this matrix decay without a significant gap.
@@
(Spectra - filter factors in fp64) - We begin with a comparison of our filter factor results presented herein by evaluating the filtering behavior of MP-IR on the Spectra problem contaminated by $1\%$ noise in fp64. For this experiment we chose the regularization parameter $\alpha^{2}=1e\text{-}2$; in the restoration part of this example we expand on this choice. Figure 1 displays the filter factors from Theorem 4.3, Theorem 4.8, and the effective filter factors compared to those of Tikhonov (24) when considering all computations in fp64 after $1$ iteration. We comment that there is no visual discernible difference amongst the plots thereby experimentally verifying that: (1) the filter factors of Theorem 4.3 and Theorem 4.8 coincide when all floating point computations are done in fp64, i.e. when $(\text{Pr}_{1},\,\text{Pr}_{2},\,\text{Pr}_{3})=(1,1,1)$ and (2) that all three sets of filter factors match those of Tikhonov. Table 2 (below) numerically confirms the agreement to unit round-off between Theorem 4.8 and the effective filter factors in fp64 working precision.
@@
Figure 1: Spectra example: filter factors of standard Tikhonov (24) compared against filter factors of iterative refinement computed using: (a) the effective filter factors (31), (b) the filter factors computed by Theorem 4.3 (23), and (c) the filter factors computed by Theorem 4.8 (30). The filter factors are compared using $\alpha^{2}=1e\text{-}2$ after $1$ iteration with $1\%$ noise.
@@
(Spectra - filter factors in mixed precision) - We utilize Theorem 4.8 to investigate the filter factors of MP-IR and compare them against the effective filter factors that come from the iterates computed by Algorithm 2 where again we utilize a noise level of $1\%$ and $\alpha^{2}=1e\text{-}2$. In Table 2 we provide the summary statistics of the absolute difference between $\Phi^{(k)}$ and $\Omega^{(k)}$ of the 64 entries computed by (31). We note that for most precision combinations investigated, that by the $5^{th}$ iteration the mean error had converged to the unit round-off of the precision associated with Pr1 (see Table 1). We illustrate in Figure 2 that compared to the behavior summarized in Figure 1 of filter factors computed in fp64 that if the precision combination is changed to $(\text{Pr}_{1},\text{Pr}_{2},\text{Pr}_{3})=(3,2,1)$ that the variation of the mean entry-wise value of $|\Phi^{(k)}-\Omega^{(k)}|$ takes on approximately the unit round-off of the precision associated with the preconditioner (i.e., Pr1). This can be observed in the variation of sequential entries in the top row of Figure 2. Furthermore, we observe in the bottom row of the figure that the absolute largest deviation between the effective filter factors and those computed by Theorem 4.8 arise in the first iteration for index numbers corresponding to a rapidly changing derivative. We surmise these differences may be a result of the orthogonality assumption made throughout or of the choice of the regularization parameter.
@@
(1,1,1) $1$ $-142.0e\text{-}$ 14 $-191.1e\text{-}$ 19 $-133.8e\text{-}$ 13 $-146.0e\text{-}$ 14
$5$ $-154.0e\text{-}$ 15 00 $-131.4e\text{-}$ 13 $-141.8e\text{-}$ 14
$10$ $-154.2e\text{-}$ 15 $-213.1e\text{-}$ 21 $-131.6e\text{-}$ 13 $-142.0e\text{-}$ 14
(2,1,1) $1$ $-52.1e\text{-}$ 5 $-111.3e\text{-}$ 11 $-43.8e\text{-}$ 4 $-56.7e\text{-}$ 5
$5$ $-62.7e\text{-}$ 6 $-142.3e\text{-}$ 14 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
$10$ $-62.7e\text{-}$ 6 $-142.3e\text{-}$ 14 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
(2,2,1) $1$ $-52.1e\text{-}$ 5 $-111.3e\text{-}$ 11 $-43.8e\text{-}$ 4 $-56.7e\text{-}$ 5
$5$ $-62.7e\text{-}$ 6 $-131.3e\text{-}$ 13 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
$10$ $-62.7e\text{-}$ 6 $-131.3e\text{-}$ 13 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
(2,2,2) $1$ $-52.0e\text{-}$ 5 $-129.8e\text{-}$ 12 $-43.8e\text{-}$ 4 $-56.4e\text{-}$ 5
$5$ $-62.6e\text{-}$ 6 $-131.6e\text{-}$ 13 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
$10$ $-62.8e\text{-}$ 6 $-146.6e\text{-}$ 14 $-55.3e\text{-}$ 5 $-68.0e\text{-}$ 6
(3,2,1) $1$ $-35.6e\text{-}$ 3 $-81.2e\text{-}$ 8 $-26.0e\text{-}$ 2 $-21.4e\text{-}$ 2
$5$ $-44.6e\text{-}$ 4 $-101.6e\text{-}$ 10 $-22.1e\text{-}$ 2 $-32.6e\text{-}$ 3
$10$ $-44.6e\text{-}$ 4 $-101.6e\text{-}$ 10 $-22.1e\text{-}$ 2 $-32.6e\text{-}$ 3
(3,2,2) $1$ $-35.6e\text{-}$ 3 $-81.2e\text{-}$ 8 $-26.0e\text{-}$ 2 $-21.4e\text{-}$ 2
$5$ $-44.6e\text{-}$ 4 $-101.6e\text{-}$ 10 $-22.1e\text{-}$ 2 $-32.6e\text{-}$ 3
$10$ $-44.6e\text{-}$ 4 $-101.6e\text{-}$ 10 $-22.1e\text{-}$ 2 $-32.6e\text{-}$ 3
(3,3,2) $1$ $-35.1e\text{-}$ 3 $-81.2e\text{-}$ 8 $-26.0e\text{-}$ 2 $-21.3e\text{-}$ 2
$5$ $-46.0e\text{-}$ 4 $-111.3e\text{-}$ 11 $-22.3e\text{-}$ 2 $-32.9e\text{-}$ 3
$10$ $-46.0e\text{-}$ 4 $-111.4e\text{-}$ 11 $-22.3e\text{-}$ 2 $-32.9e\text{-}$ 3
(3,3,1) $1$ $-35.1e\text{-}$ 3 $-81.2e\text{-}$ 8 $-26.0e\text{-}$ 2 $-21.3e\text{-}$ 2
$5$ $-46.0e\text{-}$ 4 $-111.4e\text{-}$ 11 $-22.3e\text{-}$ 2 $-32.9e\text{-}$ 3
$10$ $-46.0e\text{-}$ 4 $-111.4e\text{-}$ 11 $-22.3e\text{-}$ 2 $-32.9e\text{-}$ 3
(3,3,3) $1$ $-21.9e\text{-}$ 2 $-81.0e\text{-}$ 8 $-14.3e\text{-}$ 1 $-26.7e\text{-}$ 2
$5$ $-31.6e\text{-}$ 3 $-102.1e\text{-}$ 10 $-22.4e\text{-}$ 2 $-34.5e\text{-}$ 3
$10$ $-49.5e\text{-}$ 4 $-107.5e\text{-}$ 10 $-21.6e\text{-}$ 2 $-32.5e\text{-}$ 3
(Spectra - reconstruction) - We now compare MP-IR versus AIR on the Spectra problem. As our interests have focused on the filtering properties of MP-IR on the Tikhonov problem, we did not investigate methodologies for determining regularization parameters nor did we consider any stopping criterion. Both topics deserve their own dedicated studies in the context of mixed precision computing for inverse problems. In our numerical investigation we chose 3 magnitudes of regularization parameters: 1e-2, 1e-3, and 1e-4 as well as two noise levels: $0.5\%$ and $3\%$ to consider the broad behavior of both methods. For method comparison, we utilize the relative reconstructive error (RRE) defined by
@@
where $x$ and $x^{(k)}$ denote the true solution and $k^{th}$ approximate solution determined by MP-IR or AIR, respectively. Using the RRE, we focus on the stability of the solutions by considering what we will refer to as the stable RRE (sRRE) which is computed as the mean of the 3rd through the 10th iterates, whose iterate numbers were chosen to reflect relatively stable solution behavior. We also consider the standard deviation values for the same choice of iterates. In our numerical studies, the precision variants of MP-IR did not require more than 10 iterations to converge to an RRE value that would deviate more than the standard deviation of the sRRE.
Table 3 provides the sRRE and standard deviation values at various precisions for MP-IR as well as AIR as a function of the regularization parameter and the noise level of the problem. For succinctness in the table, we did not provide the results for $\alpha^{2}=1e\text{-}2$ as the presented choices resulted in smaller sRRE values for MP-IR when considered for both noise levels. When AIR was investigated closely, we found that it was stable over the considered 10 iterations with a sRRE of $-16.26e\text{-}$ 1 and standard deviation of $-54.0e\text{-}$ 5 for a choice of $\alpha^{2}=1e\text{-}1$ and $3\%$ noise, however, this is a significantly worse error than was attainable by the MP-IR method shown in the table indicating a superior MP-IR performance for these choices of $\alpha^{2}$.
We observed that the sRRE values for precision combinations involving fp16 for the preconditioner (i.e., $\text{Pr}_{1}=3$) could experience larger sRRE variation - this is shown graphically in Figure 3 as well as in the standard deviation columns of Table 3. Specifically, we note that while the sRRE values for all precision combinations were fairly consistent across the same level of noise, regularization parameter, and precision combination, the standard deviation of the sRRE was found to be closer to the unit round-off of Pr1. This aligns with theoretical expectations given that this is the same precision that preconditioner is computed in. This variation can be observed directly in Figure 3a for combination (3, 3, 3)333(3,\,3,\,3)( 3 , 3 , 3 ), and additionally for combinations (3, 3, 1)331(3,\,3,\,1)( 3 , 3 , 1 ) and (3, 3, 2)332(3,\,3,\,2)( 3 , 3 , 2 ) in Figure 3b. From this experiment it appears that variability in RRE can be reduced by balancing the choice of the preconditioner in lower precision with higher working and residual precisions.
Table 3: Spectra example: stable relative reconstructive error (i.e., the average RRE for iterations $3$ $10$) and the standard deviation for MP-IR and AIR methods as a function of the regularization parameter for $0.5\%$ and $3\%$ noise. MP-IR errors are provided for various mixed precision combinations.

### v3

@@ 3     $\left(\sigma_{M,j}^{2}+\alpha^{2}\right)\in\mathbb{R},\quad\text{for }j=1,2,\dots,n$ (in Pr1)
4Output: $\Psi^{(k+1)}\in\mathbb{R}^{n}$
5for $k=0,1,2,\dots$ do
@@
[…]
@@
with the $j^{th}$ filter factor per $k^{th}$ iterative step and indicated precision given by
@@
1Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2Input: $\Phi^{(0)}=0\in\mathbb{R}^{n},\,\,\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\alpha>0$
3Output: $\Phi^{(k+1)}\in\mathbb{R}^{n}$
[…]
5for $k=0,1,2,\dots$ do
[…]
@@
[…]
@@ 5.1 Effective filter factors
To assess the experimental validity of our filter factor results from Section 4 for IR on the Tikhonov problem we compare against experimentally derived filter factors which we term effective filter factors. Using (5), the effective filter factors $\omega^{(k)}_{j}$ of an approximate solution at the $k^{th}$ iteration may be computed as follows
@@ 5.2 Low precision simulation in MATLAB
To ascertain the effectiveness of mixed precision IR on the Tikhonov problem and the experimental validity of the results presented in Section 4 we utilize the software package chop introduced by Higham and Pranesh in [18]. The chop function simulates lower precision arithmetic by rounding array entries given in a MATLAB native precision (e.g., single or double) to a target precision which is stored in a higher precision with non-utilized representation bits set to zero. Chen et al. in [6] considered the application of chop for the solution of structured inverse problems and investigated associated numerical considerations of its usage.
Various target precisions are supported by the chop software including user-customizable ones. The precisions we utilize in our numerical experiments in Section 5.3 and their associated shorthand notations are provided in Table 1. We comment that we did not experience any significant experimental differences when using precisions that did not support subnormal numbers, e.g., Google’s bfloat16111https://en.wikipedia.org/wiki/Bfloat16_\_floating-point_\_format, compared to those due to the IEEE standard [^2] https://en.wikipedia.org/wiki/IEEE_\_754#\#2019 during our numerical investigations and therefore do not include them in our results.
@@
Pr
Shorthand
Name
Exponent
Bits
Mantissa
Bits
Approx. Decimal
Round-off
IEEE
Standard
11 fp64 $11$ $52$ $1.11e\text{-}16$ Yes
22 fp32 88 $23$ $5.96e\text{-}8$ Yes
33 fp16 55 $10$ $4.88e\text{-}4$ Yes
@@ 5.3 Examples and Results
[…]
@@
(Spectra - filter factors in fp64) - We begin with a comparison of our filter factor results presented herein by evaluating the filtering behavior of MP-IR on the Spectra problem contaminated by $1\%$ noise in fp64. For this experiment we chose the regularization parameter $\alpha^{2}=1e\text{-}2$; in the restoration part of this example we expand on this choice. Figure 1 displays the filter factors from Theorem 4.3, Theorem 4.8, and the effective filter factors compared to those of Tikhonov (24) when considering all computations in fp64 after $1$ iteration. We comment that there is no visual discernible difference amongst the plots thereby experimentally verifying that: (i) the filter factors of Theorem 4.3 and Theorem 4.8 coincide when all floating point computations are done in fp64, i.e., when $(\text{Pr}_{1},\,\text{Pr}_{2},\,\text{Pr}_{3})=(1,1,1)$ and (ii) that all three sets of filter factors match those of Tikhonov. Table 2 numerically confirms the agreement to unit round-off between Theorem 4.8 and the effective filter factors in fp64 working precision.
@@
Figure 1: Spectra example: filter factors of standard Tikhonov (24) compared against filter factors of iterative refinement computed using: (a) the effective filter factors (31), (b) the filter factors computed by Theorem 4.3 (23), and (c) the filter factors computed by Theorem 4.8 (30). The filter factors are compared using $\alpha^{2}=1e\text{-}2$ after 11 iteration with $1\%$ noise.
@@
(Spectra - filter factors in mixed precision) - We utilize Theorem 4.8 to investigate the filter factors of MP-IR and compare them against the effective filter factors that come from the iterates computed by Algorithm 2 where again we utilize a noise level of $1\%$ and $\alpha^{2}=1e\text{-}2$. In Table 2 we provide the summary statistics of the absolute difference between $\Phi^{(k)}$ and $\Omega^{(k)}$ of the 64 entries computed by (31). We note that for most precision combinations investigated, that by the $5^{th}$ iteration the mean error had converged to the unit round-off of the precision associated with Pr1 (see Table 1).
We illustrate in Figure 2 that compared to the behavior summarized in Figure 1 of filter factors computed in fp64 that if the precision combination is changed to $(\text{Pr}_{1},\text{Pr}_{2},\text{Pr}_{3})=(3,2,1)$ that the variation of the mean entry-wise value of $|\Phi^{(k)}-\Omega^{(k)}|$ takes on approximately the unit round-off of the precision associated with the preconditioner (i.e., Pr1). This can be observed in the variation of sequential entries in the top row of Figure 2. Furthermore, we observe in the bottom row of the figure that the absolute largest deviation between the effective filter factors and those computed by Theorem 4.8 arise in the first iteration for index numbers corresponding to a rapidly changing derivative. We surmise these differences may be a result of the orthogonality assumption made throughout or of the choice of the regularization parameter. We close by commenting that a finite precision error analysis is expected to support these results and to help explain the impact of the preconditioner on the error.
@@
(1,1,1) 11 $2.0e\text{-}14$ $1.1e\text{-}19$ $3.8e\text{-}13$ $6.0e\text{-}14$
55 $4.0e\text{-}15$ 0 $1.4e\text{-}13$ $1.8e\text{-}14$
$10$ $4.2e\text{-}15$ $3.1e\text{-}21$ $1.6e\text{-}13$ $2.0e\text{-}14$
(2,1,1) 11 $2.1e\text{-}5$ $1.3e\text{-}11$ $3.8e\text{-}4$ $6.7e\text{-}5$
55 $2.7e\text{-}6$ $2.3e\text{-}14$ $5.2e\text{-}5$ $7.8e\text{-}6$
$10$ $2.7e\text{-}6$ $2.3e\text{-}14$ $5.2e\text{-}5$ $7.8e\text{-}6$
(2,2,1) 11 $2.1e\text{-}5$ $1.3e\text{-}11$ $3.8e\text{-}4$ $6.7e\text{-}5$
55 $2.7e\text{-}6$ $1.3e\text{-}13$ $5.2e\text{-}5$ $7.8e\text{-}6$
$10$ $2.7e\text{-}6$ $1.3e\text{-}13$ $5.2e\text{-}5$ $7.8e\text{-}6$
(2,2,2) 11 $2.0e\text{-}5$ $9.8e\text{-}12$ $3.8e\text{-}4$ $6.4e\text{-}5$
55 $2.6e\text{-}6$ $1.6e\text{-}13$ $5.2e\text{-}5$ $7.8e\text{-}6$
$10$ $2.8e\text{-}6$ $6.6e\text{-}14$ $5.3e\text{-}5$ $8.0e\text{-}6$
(3,2,1) 11 $5.6e\text{-}3$ $1.2e\text{-}8$ $6.0e\text{-}2$ $1.4e\text{-}2$
55 $4.6e\text{-}4$ $1.6e\text{-}10$ $2.1e\text{-}2$ $2.6e\text{-}3$
$10$ $4.6e\text{-}4$ $1.6e\text{-}10$ $2.1e\text{-}2$ $2.6e\text{-}3$
(3,2,2) 11 $5.6e\text{-}3$ $1.2e\text{-}8$ $6.0e\text{-}2$ $1.4e\text{-}2$
55 $4.6e\text{-}4$ $1.6e\text{-}10$ $2.1e\text{-}2$ $2.6e\text{-}3$
$10$ $4.6e\text{-}4$ $1.6e\text{-}10$ $2.1e\text{-}2$ $2.6e\text{-}3$
(3,3,2) 11 $5.1e\text{-}3$ $1.2e\text{-}8$ $6.0e\text{-}2$ $1.3e\text{-}2$
55 $6.0e\text{-}4$ $1.3e\text{-}11$ $2.3e\text{-}2$ $2.9e\text{-}3$
$10$ $6.0e\text{-}4$ $1.4e\text{-}11$ $2.3e\text{-}2$ $2.9e\text{-}3$
(3,3,1) 11 $5.1e\text{-}3$ $1.2e\text{-}8$ $6.0e\text{-}2$ $1.3e\text{-}2$
55 $6.0e\text{-}4$ $1.4e\text{-}11$ $2.3e\text{-}2$ $2.9e\text{-}3$
$10$ $6.0e\text{-}4$ $1.4e\text{-}11$ $2.3e\text{-}2$ $2.9e\text{-}3$
(3,3,3) 11 $1.9e\text{-}2$ $1.0e\text{-}8$ $4.3e\text{-}1$ $6.7e\text{-}2$
55 $1.6e\text{-}3$ $2.1e\text{-}10$ $2.4e\text{-}2$ $4.5e\text{-}3$
$10$ $9.5e\text{-}4$ $7.5e\text{-}10$ $1.6e\text{-}2$ $2.5e\text{-}3$
(Spectra - reconstruction) - We now compare MP-IR versus AIR on the Spectra problem. As our interests have focused on the filtering properties of MP-IR on the Tikhonov problem, we did not investigate methodologies for determining regularization parameters, nor did we consider any stopping criterion. Both topics deserve their own dedicated studies in the context of mixed precision computing for inverse problems. In our numerical investigation we chose 3 magnitudes of regularization parameters: 1e-2, 1e-3, and 1e-4 as well as two noise levels: $0.5\%$ and $3\%$ to consider the broad behavior of both methods. For method comparison, we utilize the relative reconstructive error (RRE) defined by
@@
where $x$ and $x^{(k)}$ denote the true solution and $k^{th}$ approximate solution determined by MP-IR or AIR, respectively. Using the RRE, we focus on the stability of the solutions by considering what we will refer to as the stable RRE (sRRE) which is computed as the mean of the 3rd through the 10th iterates, whose iterate numbers were chosen to reflect relatively stable solution behavior. We also consider the standard deviation values for the same choice of iterates. In our numerical studies, the variants of MP-IR did not require more than 10 iterations to converge to an RRE value that would deviate more than a standard deviation from the sRRE.
Table 3 provides the sRRE and standard deviation values at various precisions for MP-IR as well as AIR as a function of the regularization parameter and the noise level of the problem. For succinctness in the table, we did not provide the results for $\alpha^{2}=1e\text{-}2$ as the presented choices resulted in smaller sRRE values for MP-IR when considered for both noise levels. When AIR was investigated closely, we found that it was stable over the considered 10 iterations with a sRRE of $6.26e\text{-}1$ and standard deviation of $4.0e\text{-}5$ for a choice of $\alpha^{2}=1e\text{-}1$ and $3\%$ noise, however, this is a significantly worse error than was attainable by the MP-IR method shown in the table indicating a superior MP-IR performance for these choices of $\alpha^{2}$.
We observed that the sRRE values for precision combinations involving fp16 for the preconditioner (i.e., $\text{Pr}_{1}=3$) could experience larger sRRE variation - this is shown graphically in Figure 3 as well as in the standard deviation columns of Table 3. Specifically, we note that while the sRRE values for all precision combinations were fairly consistent across the same level of noise, regularization parameter, and precision combination, the standard deviation of the sRRE was found to be closer to the unit round-off of Pr1. This aligns with theoretical expectations given that this is the same precision that preconditioner is computed in. This variation can be observed directly in Figure 3a for combination (3, 3, 3)(3,\,3,\,3), and additionally for combinations (3, 3, 1)(3,\,3,\,1) and (3, 3, 2)(3,\,3,\,2) in Figure 3b. From this experiment it appears that variability in RRE can be reduced by balancing the choice of the preconditioner in lower precision with higher working and residual precisions.
Table 3: Spectra example: stable relative reconstructive error (i.e., the average RRE for iterations 33 - $10$) and the standard deviation for MP-IR and AIR methods as a function of the regularization parameter for $0.5\%$ and $3\%$ noise. MP-IR errors are provided for various mixed precision combinations.